An astronaut in his space suit and with a propulsion unit (empty of its gas propellant) strapped to his back has a mass of 146 kg. The astronaut begins a space walk at rest, with a completely filled propulsion unit. During the space walk, the unit ejects some gas with a velocity of 32 m/s. as a result, the astronaut recoils with a velocity of .39 m/s. After the gas is ejected, the mass of the astronaut (now wearing a partially empty propulsion unit) is 165 kg. what percentage of the gas was ejected from the completely filled propulsion unit?
Answer #1

Mass of left out unit = 165 - 146 = 19 kg

Mass of ejected gas, m = (165 x 0.39)/32 = 2.01 kg

Total mass = 19 + 2.01 = 21.01 kg

% increase = (2.01/21.01) x 100 = 9.57 %

Comment in case any doubt please rate my answer .....